OpenStudy (anonymous):

HELP PLEASE!!! Find the volume of a cone with a diameter of 5 inches and a height that is 3 times the radius. Use 3.14 for pi, and round your answer to the nearest hundredth. 49.06 in.3 7851 in.3 147.19 in.3 392.50 in.3 @campbell_st @One098

OpenStudy (anonymous):

@iGreen @AnswerMyQuestions

OpenStudy (anonymous):

@Directrix please help me!!!!

OpenStudy (campbell_st):

ok, so the diameter is 5 inches, what is the radius? (2 x radius = 1 x diameter)

OpenStudy (campbell_st):

the volume formula is \[V = \frac{1}{3} \pi r^2h\] so when you find the radius (r) remember the height (h) = 3r when you get both values, and you as asked to use pi = 3.14 the substitute the values into \[V = \frac{1}{3} \times 3.14 \times r^2 \times h\] hope it helps